# Generalized Bayesian Multidimensional Scaling and Model Comparison

### v2

@@ Abstract
Multidimensional scaling (MDS) is widely used to reconstruct a low-dimensional representation of high-dimensional data while preserving pairwise distances. However, Bayesian MDS approaches based on Markov chain Monte Carlo (MCMC) face challenges in model generalization and comparison. To address these limitations, we propose a generalized Bayesian multidimensional scaling (GBMDS) framework that accommodates non-Gaussian errors and diverse dissimilarity metrics for improved robustness. We develop an adaptive annealed Sequential Monte Carlo (ASMC) algorithm for Bayesian inference, leveraging an annealing schedule to enhance posterior exploration and computational efficiency. The ASMC algorithm also provides a nearly unbiased marginal likelihood estimator, enabling principled Bayesian model comparison across different error distributions, dissimilarity metrics, and dimensional choices. Using synthetic and real data, we demonstrate the effectiveness of the proposed approach. Our results show that ASMC-based GBMDS achieves superior computational efficiency and robustness compared to MCMC-based methods under the same computational budget. The implementation of our proposed method and applications are available at https://github.com/SFU-Stat-ML/GBMDS.
Sequential Monte Carlo,
dimension reduction,
adaptive inference,
robustness,
skewness,
visualization,
@@
volume: TBAissue: TBA
[…]
, and
@@ 1 Introduction
MDS methods are broadly categorized into metric and non-metric approaches. Metric MDS assumes that dissimilarities represent numerical distances and satisfy metric properties such as symmetry and the triangle inequality. It is particularly appropriate when the underlying data structure follows Euclidean geometry or when a specific distance function, such as the Euclidean or Mahalanobis distance, is justified. Non-metric MDS, on the other hand, is designed for ordinal or rank-based dissimilarities, where the absolute values of distances are not meaningful, and only the relative ordering of dissimilarities matters. This method is widely used in applications where dissimilarities do not adhere to strict metric properties, such as in psychological or perceptual studies. Despite differences in assumptions, both MDS approaches aim to embed high-dimensional data into a lower-dimensional space while preserving the original similarity structure as faithfully as possible. Our study primarily focuses on metric MDS. For a comprehensive review of modern MDS methods, see Borg and Groenen (2005).
Classical multidimensional scaling (CMDS), introduced by Torgerson (1952), is a foundational method for metric MDS that reconstructs a low-dimensional Euclidean representation from a given dissimilarity matrix. CMDS is effective when the given pairwise dissimilarities are precisely equal to the Euclidean distances and when the optimal low-dimensional configuration is accurately specified (Oh and Raftery, 2001). While CMDS is computationally efficient and widely used, its performance is limited when these assumptions do not hold. In many practical applications, dissimilarities may contain noise or measurement errors, leading to distortions in the recovered configuration. In such cases, alternative approaches, such as non-metric MDS or Bayesian formulations, may provide more robust solutions.
@@
Oh and Raftery (2001) introduced a Bayesian multidimensional scaling (BMDS) approach that models observed dissimilarities as noisy observations of underlying Euclidean distances, incorporating measurement errors explicitly within a probabilistic model. Object locations in the low-dimensional space are estimated using Markov chain Monte Carlo (MCMC) sampling. Their results indicate that BMDS provides more accurate embeddings than CMDS, particularly when dissimilarities contain substantial measurement errors, the Euclidean assumption is violated, or the latent dimension is misspecified.
@@
The Bayesian approach to the MDS problem has gained increasing attention due to its ability to incorporate prior information and handle measurement uncertainty. Several extensions of BMDS have been proposed to enhance its applicability. Oh and Raftery (2007) integrated Bayesian model-based clustering within the BMDS framework in Oh and Raftery (2001) to facilitate clustering in high-dimensional spaces. Bakker and Poole (2013) modeled observed distances using a log-normal distribution and estimated posteriors via squared error loss minimization. Lin and Fong (2019) employed a $t$-distribution for object locations, improving robustness against outliers and enabling variable selection through a latent multivariate regression structure. Liu et al. (2024) extended BMDS to hyperbolic manifolds, where dissimilarities are modeled as hyperbolic distances. Advancements in sampling methods have further improved the efficiency of BMDS. Gronau and Lee (2020) introduced differential evolution MCMC to enhance sampling performance, particularly in the context of psychologically interpretable metrics such as Euclidean and Minkowski distances. Holbrook et al. (2020) leveraged Hamiltonian Monte Carlo (HMC) (Neal et al., 2011) and massive parallelization using multi-core CPUs and GPUs to accelerate inference in BMDS, with applications in phylogenetics. Despite these developments, a comprehensive Bayesian framework that accommodates non-Gaussian errors and extends beyond Euclidean spaces remains an open challenge. Moreover, while MCMC-based inference is widely used, its computational limitations motivate the exploration of alternative methods, such as sequential Monte Carlo (SMC), to improve scalability and efficiency.
Sequential Monte Carlo (SMC) methods have becoming an important alternative method for conducting Bayesian inference (see Doucet et al., 2001; Doucet and Johansen, 2009; Chopin and Papaspiliopoulos, 2020, for an introduction to SMC). In general, SMC employs a population of weighted particles that evolve sequentially, allowing for greater parallelization and adaptability. SMC methods approximate a sequence of probability distributions using importance sampling and resampling mechanisms, yielding a flexible framework for unbiased estimation of marginal likelihoods (Doucet et al., 2006). A particularly relevant variant of SMC is the SMC sampler (Del Moral et al., 2006; Dai et al., 2022), also known as annealed SMC (ASMC) (Wang et al., 2021). ASMC leverages an annealing schedule, similar to simulated tempering, to smoothly transition from an easy-to-sample distribution to the target posterior. This gradual adaptation reduces the risk of getting trapped in local modes and improves exploration of complex posterior landscapes under the same computational budget as standard MCMC. Notably, ASMC yields unbiased marginal likelihood estimators when using a fixed sequence of annealing parameters and nearly unbiased estimators when employing an adaptive annealing scheme. The estimated marginal likelihood enables robust Bayesian model comparison via Bayes factors (Jeffreys, 1935; Han and Carlin, 2001; Zhou et al., 2016; Wang et al., 2020). The efficiency of ASMC has been demonstrated in diverse applications, including epidemiology (Del Moral et al., 2012), phylogenetics (Wang et al., 2020), and solving nonlinear differential equation systems (Wang et al., 2021). For a comprehensive review of recent advances in ASMC, see Dai et al. (2022).
@@
The existing BMDS methods have several limitations. First, most BMDS methods assume Euclidean distance as the primary dissimilarity metric. However, Euclidean distance may not be suitable for certain applications. In medical imaging and 3D face recognition, surface alignment and shape comparison often rely on the Gromov-Hausdorff distance (Mémoli, 2011) or the partial embedding distance (Bronstein et al., 2006), which capture geometric distortions more effectively than Euclidean metrics. In text mining, Cosine dissimilarity (Li and Han, 2013) is preferred, as it measures the angle between document vectors and remains invariant to document length. Similarly, Jaccard dissimilarity (Jaccard, 1901) is widely used for set-based comparisons, particularly in short-text clustering. The lack of BMDS frameworks accommodating these alternative dissimilarities limits its applicability across diverse fields. Second, existing BMDS methods predominantly assume Gaussian noise in the observed dissimilarities, resulting in a lack of robustness and generality. Real-world datasets often exhibit heavy-tailed or skewed error distributions, leading to biased or inefficient estimates when using Gaussian-based models. Developing BMDS methods that incorporate more flexible error structures, such as Student’s $t$-distributed errors or nonparametric approaches, remains an open problem. Third, research on model comparison within the BMDS framework remains scarce, particularly concerning the impact of different dissimilarity distributions. Most existing work focuses on comparing Bayesian and frequentist MDS solutions in specific application domains. A systematic evaluation of BMDS models incorporating diverse dissimilarity metrics and error structures would provide deeper insights into their relative performance. Fourth, the increasing availability of high-dimensional data presents computational challenges for BMDS. Bayesian inference, while offering flexibility and uncertainty quantification, is often computationally expensive. Scaling BMDS methods to handle large datasets efficiently remains a critical issue, necessitating advances in inference techniques such as variational approximations, parallelized MCMC, or SMC methods.
@@
Our contributions can be summarized as follows: (i) We generalize BMDS to accommodate non-Gaussian errors in the pairwise dissimilarities. This allows our model to handle data with heavy-tailed or skewed distributions, improving robustness and accuracy. (ii) Unlike conventional BMDS models restricted to Euclidean distances, GBMDS supports a broad class of dissimilarity measures, making it applicable to diverse domains, such as text mining. (iii) We introduce an efficient adaptive Bayesian inference framework based on annealed SMC. This approach mitigates computational burdens in large-scale applications while leveraging existing Metropolis-Hastings proposals, ensuring ease of implementation. (iv) Our framework provides nearly unbiased estimators of marginal likelihoods as a byproduct of sampling, which facilitates model selection through Bayes factors. This enables rigorous comparison between competing BMDS models with different error structures, dissimilarity metrics, and dimensions. (v) We evaluate the proposed GBMDS framework in three simulation studies and four real-world applications, demonstrating superior estimation accuracy and robustness across diverse dissimilarity metrics compared to benchmark methods.
@@ 2 BMDS Models
Let $\mathbf{x}_{i}=(x_{i,1},\ldots,x_{i,p})^{\top}\in\mathbb{R}^{p}$ be the unobserved vector representing the values of $p$ significant attributes in object $i$. The goal of MDS methods is to find the set of points $\textbf{X}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$ such that $d_{i,j}$ and $\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{p}$ are as close as possible, where $\|\cdot\|_{p}$ represents the $L^{p}$ norm. In such a manner, the given dissimilarities are well-reproduced by the resulting configuration. We refer to this process as object configuration (Oh and Raftery, 2001), which describes the estimation of values for objects’ significant attributes.
CMDS is a commonly used dimension reduction technique for metric MDS developed by Torgerson (1952). CMDS assumes the dissimilarity to be Euclidean and takes the pairwise dissimilarities as inputs and outputs the coordinates of points in a low-dimensional space up to locations, rotations and reflections. Numerical optimization techniques can be used to find a solution to the minimization problem below:
@@ 2.1 Generalized Bayesian multidimensional scaling
Compared with the BMDS framework proposed in Oh and Raftery (2001), we do not restrict $d_{i,j}$ to be accompanied by Gaussian errors. The previous BMDS framework may be inadequate when dealing with dissimilarity measures that are subject to random errors or those that are non-Euclidean in nature. In addition, the assumption of utilizing a truncated Gaussian distribution to model the errors is inadequate in the presence of outliers. The presence of outliers can lead to increased uncertainty surrounding unobserved dissimilarities ($\delta_{i,j}$’s) beyond what can be accounted for by the tails of Gaussian distributions. We will refer to the framework in Oh and Raftery (2001) as the standard BMDS throughout this paper.
@@ 2.1.2 GBMDS with truncated skewed Gaussian distribution
For a given matrix of dissimilarities D, the likelihood function, $l$, of the latent variables $\textbf{X}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$, unknown parameters $\sigma^{2}$, $\psi$ and $U$ under $\mathcal{M}_{TSN}$, can be written as:
$l\left(\textbf{D}|\textbf{X},\sigma^{2},\psi,U,\mathcal{M}_{TSN}\right)$ $\propto\left(\sigma^{2}\right)^{-\frac{m}{2}}\prod_{\begin{subarray}{c}i,j=1\\ i>j\end{subarray}}^{n}\left\{F_{\delta_{i,j},\sigma,\psi}\left(U\right)-F_{\delta_{i,j},\sigma,\psi}\left(0\right)\right\}^{-1}$
×$\times\exp\left\{-\frac{1}{2\sigma^{2}}SSR\right\}\times\prod_{\begin{subarray}{c}i,j=1\\ i>j\end{subarray}}^{n}\Phi\left(\psi\frac{d_{i,j}-\delta_{i,j}}{\sigma}\right),$ (4)
where $F_{\delta_{i,j},\sigma,\psi}(\cdot)$ is the cdf of skewed Gaussian distribution with location $\delta_{i,j}$, scale $\sigma$, and shape $\psi$, $SSR=\sum_{i>j}\left(d_{i,j}-\delta_{i,j}\right)^{2}$ is the sum of squared residuals, $\Phi(\cdot)$ is the standard Gaussian cdf, and $m=n(n-1)/2$ is the total number of dissimilarities for $n$ objects.

### v3

@@ Abstract
Multidimensional scaling (MDS) is a fundamental dimension reduction technique widely applied across diverse fields to represent objects as points in a lower-dimensional space based on pairwise dissimilarities. While classical and scalable non-Bayesian MDS methods are valued for their computational efficiency, they primarily focus on optimization, yielding only point estimates and thus neglecting model uncertainty. This limitation motivates the development of Bayesian MDS (BMDS) frameworks, which offer a probabilistic perspective, model uncertainty, and provide posterior distributions for principled uncertainty quantification. However, existing BMDS methods predominantly assume Euclidean distance and Gaussian noise in the observed dissimilarities, which limits robustness and applicability across diverse domains like text mining. Furthermore, Bayesian inference for these models faces computational challenges when scaling to large datasets. To overcome these issues, we propose a Generalized Bayesian Multidimensional Scaling (GBMDS) framework that incorporates flexible dissimilarity metrics and robust non-Gaussian error structures, prioritizing uncertainty quantification, robustness, and model flexibility. To facilitate efficient inference and robust model selection within this framework, we design an adaptive annealed Sequential Monte Carlo (ASMC) algorithm. This ASMC approach mitigates computational burdens in large-scale applications, leverages existing MCMC proposals for ease of implementation, and provides nearly unbiased estimators of marginal likelihoods, enabling rigorous Bayesian model comparison via Bayes factors. The GBMDS framework thus enhances estimation accuracy and robustness by accommodating complex dissimilarities and heavy-tailed distributions.
@@
Sequential Monte Carlo, dimension reduction, adaptive inference, robustness, skewness, visualization
@@ 1 Introduction
MDS methods are broadly categorized into metric and non-metric approaches. Metric MDS assumes that dissimilarities represent numerical distances and satisfy metric properties such as symmetry and the triangle inequality. It is particularly appropriate when the underlying data structure follows Euclidean geometry or when a specific distance function, such as the Euclidean or Mahalanobis distance, is justified. Non-metric MDS, on the other hand, is designed for ordinal or rank-based dissimilarities, where the absolute values of distances are not meaningful, and only the relative ordering of dissimilarities matters. This method is widely used in applications where dissimilarities do not adhere to strict metric properties, such as in psychological or perceptual studies. Despite differences in assumptions, both MDS approaches aim to embed high-dimensional data into a lower-dimensional space while preserving the original similarity structure as faithfully as possible. Our study primarily focuses on metric MDS. For a comprehensive review of modern MDS methods, see [borg2005modern].
Classical multidimensional scaling (CMDS), introduced by torgerson1952multidimensional, is an essential method for metric MDS that reconstructs a low-dimensional Euclidean representation from a given dissimilarity matrix. CMDS is computationally efficient for small datasets; however, its scalability is limited by the extensive memory and time requirements associated with storing and processing the full dissimilarity matrix for large sample sizes. To address this, scalable alternatives have been developed. For example, landmark MDS algorithm [de2004sparse] first applies CMDS to a representative subset of data (landmark points) and then projects the remaining data onto the resulting low-dimensional configuration using a distance-based triangulation procedure. Another popular approach includes Divide-and-Conquer MDS [delicado2025multidimensional], which preserves local structure by splitting the dataset into overlapping patches, performing independent MDS on each, and then aligning and combining the partial configurations via Procrustes transformation. For a comprehensive review of the scalable non-Bayesian MDS methods, see delicado2025multidimensional.
While classical and scalable non-Bayesian MDS methods are valued for their computational efficiency, they primarily focus on optimization and yield only point estimates. This limitation motivates the development of Bayesian MDS frameworks, which offer a probabilistic perspective on dimension reduction. The Bayesian approach explicitly models uncertainty and noise in the observed dissimilarities and provides posterior distributions over the latent coordinates, allowing for principled uncertainty quantification and greater flexibility through the incorporation of prior structures. Given these methodological advantages, and since we prioritize uncertainty quantification, robustness, and model flexibility over scalability, our paper will focus on Bayesian formulations of MDS, leaving comparisons with large-scale non-Bayesian MDS methods beyond the scope of this study.
@@
oh2001bayesian introduced a Bayesian multidimensional scaling approach (BMDS) that models observed dissimilarities as noisy observations of underlying Euclidean distances, incorporating measurement errors explicitly within a probabilistic model. Object locations in the low-dimensional space are estimated using Markov chain Monte Carlo (MCMC) sampling. Their results indicate that BMDS provides more accurate embeddings than CMDS, particularly when dissimilarities contain substantial measurement errors, the Euclidean assumption is violated, or the latent dimension is misspecified.
@@
The Bayesian approach to the MDS problem has gained increasing attention due to its ability to incorporate prior information and handle measurement uncertainty. Several extensions of BMDS have been proposed to enhance its applicability. oh2007model integrated Bayesian model-based clustering within the BMDS framework in oh2001bayesian to facilitate clustering in high-dimensional spaces. bakker2013bayesian modeled observed distances using a log-normal distribution and estimated posteriors via squared error loss minimization. lin2019bayesian employed a $t$-distribution for object locations, improving robustness against outliers and enabling variable selection through a latent multivariate regression structure. liu2024bayesian extended BMDS to hyperbolic manifolds, where dissimilarities are modeled as hyperbolic distances. Advancements in sampling methods have further improved the efficiency of BMDS. gronau2020bayesian introduced differential evolution MCMC to enhance sampling performance, particularly in the context of psychologically interpretable metrics such as Euclidean and Minkowski distances. holbrook2020massive leveraged Hamiltonian Monte Carlo (HMC) [neal2011mcmc] and massive parallelization using multi-core CPUs and GPUs to accelerate inference in BMDS, with applications in phylogenetics. Despite these developments, a comprehensive Bayesian framework that accommodates non-Gaussian errors and extends beyond Euclidean spaces remains an open challenge. Furthermore, traditional MCMC-based inference often struggles with the multimodal posterior landscapes common in MDS, limiting its scalability and efficiency.
@@
The existing BMDS methods have several limitations. First, most BMDS methods assume Euclidean distance as the primary dissimilarity metric. However, Euclidean distance may not be suitable for certain applications. In medical imaging and 3D face recognition, surface alignment and shape comparison often rely on the Gromov-Hausdorff distance [memoli2011gromov] or the partial embedding distance [bronstein2006generalized], which capture geometric distortions more effectively than Euclidean metrics. In text mining, Cosine dissimilarity [li2013distance] is preferred, as it measures the angle between document vectors and remains invariant to document length. Similarly, Jaccard dissimilarity [jaccard1901etude] is widely used for set-based comparisons, particularly in short-text clustering. The lack of BMDS frameworks accommodating these alternative dissimilarities limits its applicability across diverse fields. Second, existing BMDS methods predominantly assume Gaussian noise in the observed dissimilarities, resulting in a lack of robustness and generality. Real-world datasets often exhibit heavy-tailed or skewed error distributions, leading to biased or inefficient estimates when using Gaussian-based models. Developing BMDS methods that incorporate more flexible error structures, such as Student’s $t$-distributed errors or nonparametric approaches, remains an open problem. Third, research on model comparison within the BMDS framework remains scarce, particularly concerning the impact of different dissimilarity distributions. Most existing work focuses on comparing Bayesian and frequentist MDS solutions in specific application domains. A systematic evaluation of BMDS models incorporating diverse dissimilarity metrics and error structures would provide deeper insights into their relative performance. Fourth, the increasing availability of high-dimensional data presents computational challenges for BMDS. Bayesian inference, while offering flexibility and uncertainty quantification, is often computationally expensive. Scaling BMDS methods to handle large datasets efficiently remains a critical issue, necessitating advances in inference techniques such as variational approximations, parallelized MCMC, or SMC methods.
Sequential Monte Carlo (SMC) methods have become an important alternative method for conducting Bayesian inference [see doucet2001sequential, doucet2009tutorial, chopin2020introduction, for an introduction to SMC]. In general, SMC employs a population of weighted particles that evolve sequentially, allowing for greater parallelization and adaptability. SMC methods approximate a sequence of probability distributions using importance sampling and resampling mechanisms, yielding a flexible framework for unbiased estimation of marginal likelihoods [doucet2006efficient]. A particularly relevant variant of SMC is the SMC sampler [del2006sequential, dai2022invitation], also known as annealed SMC (ASMC) [wang2021adaptive]. ASMC leverages an annealing schedule, similar to simulated tempering, to smoothly transition from an easy-to-sample distribution to the target posterior. This gradual adaptation reduces the risk of getting trapped in local modes and improves exploration of complex posterior landscapes under the same computational budget as standard MCMC. Notably, ASMC yields unbiased marginal likelihood estimators when using a fixed sequence of annealing parameters and nearly unbiased estimators when employing an adaptive annealing scheme. The estimated marginal likelihood enables robust Bayesian model comparison via Bayes factors [jeffreys1935some, han2001markov, zhou2016toward, wang2020annealed]. The efficiency of ASMC has been demonstrated in diverse applications, including epidemiology [del2012adaptive], phylogenetics [wang2020annealed], and solving nonlinear differential equation systems [wang2021adaptive]. For a comprehensive review of recent advances in ASMC, see dai2022invitation.
@@
Our contributions can be summarized as follows: (i) We generalize BMDS to accommodate non-Gaussian errors in the pairwise dissimilarities. This allows our model to handle data with heavy-tailed or skewed distributions, improving robustness and accuracy. (ii) Unlike conventional BMDS models restricted to Euclidean distances, GBMDS supports a broad class of dissimilarity measures, making it applicable to diverse domains, such as text mining. (iii) We introduce an efficient adaptive Bayesian inference framework based on annealed SMC. This approach mitigates computational burdens in large-scale applications while leveraging existing Metropolis-Hastings proposals, ensuring ease of implementation. (iv) Our framework provides nearly unbiased estimators of marginal likelihoods as a byproduct of sampling, which facilitates model selection through Bayes factors. This enables rigorous comparison between competing BMDS models with different error structures, dissimilarity metrics, and dimensions. (v) We evaluate the proposed GBMDS framework in three simulation studies and four real-world applications, demonstrating superior estimation accuracy and robustness across diverse dissimilarity metrics compared to benchmark methods. The implementation of our proposed method and applications are available at https://github.com/SFU-Stat-ML/GBMDS. The implementation of the simulation studies is available at https://github.com/nunujiarui/GBMDS_simulation.
@@ 2 BMDS Models
Let $\mathbf{x}_{i}=(x_{i,1},\ldots,x_{i,p})^{\top}\in\mathbb{R}^{p}$ be the unobserved vector representing the values of $p$ significant attributes in object $i$. The goal of MDS methods is to find the set of points $\textbf{X}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$ such that $d_{i,j}$ and $\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{p}$ are as close as possible, where $\|\cdot\|_{p}$ represents the $L^{p}$ norm. In such a manner, the given dissimilarities are well-reproduced by the resulting configuration. We refer to this process as object configuration [oh2001bayesian], which describes the estimation of values for objects’ significant attributes.
CMDS is a commonly used dimension reduction technique for metric MDS developed by torgerson1952multidimensional. CMDS assumes the dissimilarity to be Euclidean and takes the pairwise dissimilarities as inputs and outputs the coordinates of points in a low-dimensional space up to locations, rotations and reflections. Numerical optimization techniques can be used to find a solution to the minimization problem below:
@@ 2.1 Generalized Bayesian multidimensional scaling
Compared with the BMDS framework proposed in oh2001bayesian, we do not restrict $d_{i,j}$ to be accompanied by Gaussian errors. The previous BMDS framework may be inadequate when dealing with dissimilarity measures that are subject to random errors or those that are non-Euclidean in nature. In addition, the assumption of utilizing a truncated Gaussian distribution to model the errors is inadequate in the presence of outliers. The presence of outliers can lead to increased uncertainty surrounding unobserved dissimilarities ($\delta_{i,j}$’s) beyond what can be accounted for by the tails of Gaussian distributions. We will refer to the framework in oh2001bayesian as the standard BMDS throughout this paper.
@@ 2.1.2 GBMDS with truncated skewed Gaussian distribution
Let the latent variables $\textbf{X}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$. Under $\mathcal{M}_{TSN}$, for a given matrix of dissimilarities D, the likelihood function $L$ with respect to X, $\sigma^{2}$, and $\psi$, can be written as:
$L\left(\textbf{X},\sigma^{2},\psi;\textbf{D},\mathcal{M}_{TSN}\right)$ $\propto\left(\sigma^{2}\right)^{-\frac{m}{2}}\prod_{\begin{subarray}{c}i,j=1\\ i>j\end{subarray}}^{n}\left\{F_{\delta_{i,j},\sigma,\psi}\left(U\right)-F_{\delta_{i,j},\sigma,\psi}\left(0\right)\right\}^{-1}$
$\times\exp\left\{-\frac{1}{2\sigma^{2}}SSR\right\}\times\prod_{\begin{subarray}{c}i,j=1\\ i>j\end{subarray}}^{n}\Phi\left(\psi\frac{d_{i,j}-\delta_{i,j}}{\sigma}\right),$ (4)
where $F_{\delta_{i,j},\sigma,\psi}(\cdot)$ is the cdf of skewed Gaussian distribution with location $\delta_{i,j}$, scale $\sigma$, and shape $\psi$, $SSR=\sum_{i>j}\left(d_{i,j}-\delta_{i,j}\right)^{2}$ is the sum of squared residuals, $\Phi(\cdot)$ is the standard Gaussian cdf, and m=n(n1)/2m=n(n-1)/2 is the total number of dissimilarities for $n$ objects.